OpenStudy (geometry_hater):

Find each coefficent described. Coeffecient of y^3 in expression of (2y+1)^6

OpenStudy (mathmate):

The general expansion of (a+b)^n is given by C(n,0)a^n + C(n,1)a^(n-1)b + C(n,2)a^(n-2)b^2 +... C(n,n)b^n where C(n,r) is the "n choose r" function, C(n,r)=n!/(r!(n-r)!) In the given case, to get the coefficient of y^3, a=2y b=1, n=6, r=3 C(6,3)(2y)^3(1)^(6-3) =6!/(3!3!)2^3 y^3 1^3 =20*8y^3 =160y^3

OpenStudy (geometry_hater):

Thanks though this is kinda confusing to me.